OpenStudy (anonymous):

explain in complete sentences how u would divide the following expression: ^11sqrtx^5 / ^8sqrtx^3 write the simpllified answer in radical form

OpenStudy (abb0t):

\[\sqrt[n]{a^x} = a^\frac{ n }{ x }\] also know that: \[\frac{ a^n }{ a^C }= a^{n-C}\]